OpenStudy (anonymous):

Are any of these equations for the line of best fit represent data sets that have a positive linear relationship? y= -0.0025 - 0.8x y= 88.5 - 8.8x y= 13.89 - 2.9x

OpenStudy (inkyvoyd):

Well, what do you think positive linear relationship means?

OpenStudy (anonymous):

that the linear relationshp is positive :P

OpenStudy (inkyvoyd):

Here's what it means: When x increases, y increases.

OpenStudy (anonymous):

cant i do like y=mx+b and if it is positive it would be a positive relationship

OpenStudy (anonymous):

nonne have positive relationship...because the slope of all the lines are negative

OpenStudy (inkyvoyd):

@anupam1031 , can you not answer snipe?

OpenStudy (inkyvoyd):

@valeria_fl , that's not exactly how it works.

OpenStudy (anonymous):

ohh., okay, it was just a guess anyway.

OpenStudy (inkyvoyd):

Positive means that when x is increasing, y is increasing. It doesn't have to do with the specific values of x and y, it has to do with *the relationship between x and y*

OpenStudy (anonymous):

gotchya.. makes more since like that,,, ositive is positive an negative is negative.

OpenStudy (anonymous):

positive*

OpenStudy (inkyvoyd):

y= -0.0025 - 0.8x When x=0, y=-0.0025 right? When x=1, y=0.0025-0.8 right? When x increases, does y increase or decrease?

OpenStudy (inkyvoyd):

*When x=1, y=-0.0025-0.8 right?

OpenStudy (anonymous):

indeed

OpenStudy (inkyvoyd):

So, when x increases, does y increase or decrease?

OpenStudy (anonymous):

decrease

OpenStudy (inkyvoyd):

So, is it a positive or negative relationship?

OpenStudy (anonymous):

negative

OpenStudy (inkyvoyd):

Good.

OpenStudy (inkyvoyd):

y= 88.5 - 8.8x What about this one?

OpenStudy (anonymous):

Ahhh YEAHHH! :) thank you!

OpenStudy (inkyvoyd):

:P You do the same thing for every single one, and see what you get.